OpenStudy (anonymous):

Denny, Jamal, and Bobby are saving up quarters. Denny has one fifth as many as Bobby, and Jamal has three fewer than twice as many as Denny. All three of them combined have 61 quarters. Find how many each has.

OpenStudy (nikkibxox):

Lets make an equation to solve the question. Use Variables: Denny (d), Bobby (b), and Jamal (j). Use X for the unknown So we know that Denny has 1/5 or .20 as many quarters as Bobby ( 2 out of every 10) Jamal has 3 less than twice as many as D J=(D*2) -3 All combined to 61 If Denny has 2 out of every 10 quaters he would have 6 out of every 60. Denny has 6 quarters. Now that you have found out D=6 "Plug and Chug" J=(6*3)-3 J=18-3 Jamal has 15 quarters Find out how many bobby has tale 15+6 and subtract it from 61 to get 40. Overall Denny has 6 quarters Jamal has 15 and Bobby has 40

OpenStudy (anonymous):

Thank you NikkiBxox!!

OpenStudy (anonymous):

the above is incorrect... Jamal has 3 fewer than twice as many as Denny, which according to that answer would be 9 for Jamal (6 * 2 - 3). Bobby does have 40. 1/5 of that is 8, so Denny has 8. Jamal has (8 * 2 - 3) = 13. Total 13 + 8 + 40 = 61.